OpenStudy (anonymous):

Evaluate 15/J when 5=j .

OpenStudy (anonymous):

3

OpenStudy (anonymous):

15/j j=5 15/5=3

OpenStudy (ash2326):

We hav \[ \frac{15}{j}\] j=5 \[\frac{15}{5}\] we get 3